please with chi chart thank you
An anatomy teacher hypothesizes that the final grades in her class are distributed as 10% As, 23% Bs, 45% Cs, 14% Ds, and
Answer #1

6)

null hypothesis:Ho:distribution Grades distribution in sample is same as distribution in population.
Alternate hypothesis:Ha:distribution Grades distribution in sample is not same as distribution in population.
degree of freedom =categories-1= 4
for 0.05 level and 4 df :crtiical value X2 = 9.488
Decision rule: reject Ho if value of test statistic X2>9.488
applying chi square goodness of fit test:
           relative observed Expected residual Chi square
category frequency(p) Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
A 0.100 6.000 5.400 0.26 0.067
B 0.230 14.000 12.420 0.45 0.201
C 0.450 22.000 24.300 -0.47 0.218
D 0.140 8.000 7.560 0.16 0.026
F 0.080 4.000 4.320 -0.15 0.024
total 1.000 54 54 0.5347
test statistic X2 = 0.535
since test statistic does not falls in rejection region we fail to reject null hypothesis
we do not have have sufficient evidence to conclude that distribution of grades differ from claimed by teacher